### v1

@@ A.2 ClinicalAgent Algorithm
1:Input: Chief Complaint $CC$, Medical History $H$, physical examination $V$, Priori Knowledge $PK$, Number of departments arranged $K$, Number of Clinicians $N$
2:Output: Final Diagnosis $FD$, Treatment Plan $TP$
@@
5:$CT\leftarrow\emptyset$      % Clinician Team
[…]
7:for $k=1$ to $K$ do
8:     for $n=1$ to $N$ do
9:         $CL_{LLM}^{Dept}\leftarrow\text{Mapping}(DE_{k},PK,n)$      % Retrieve and call clinicians
10:         $CT.\text{append}(CL_{LLM}^{Dept})$
@@
15:$CL_{Chair}\leftarrow\text{SelectLeadClinician}(CT)$
16:$PD,LE,IE\leftarrow\emptyset,\emptyset,\emptyset$
17:for $CL_{i}$ in $CT$ do
18:     $PD_{i},LE_{i},IE_{i}\leftarrow CL_{i}(CC,H,V)$
19:     $PD.\text{append}(PD_{i});\leavevmode\nobreak\ \leavevmode\nobreak\ LE.\text{append}(LE_{i});\leavevmode\nobreak\ \leavevmode\nobreak\ IE.\text{append}(IE_{i})$
@@
21:$PD,LE,IE\leftarrow\text{Aggregate}(PD,CL_{Chair}),\text{Aggregate}(LE,CL_{Chair}),\text{Aggregate}(IE,CL_{Chair})$
@@
24:$A\leftarrow\emptyset$
25:for $test_{j}$ in $LE$ do
26:     $A_{j}\leftarrow\text{Biochemist}(test_{j},PA)$
27:     $A.\text{append}(A_{j})$
28:end for
29:
30:Entry: Radiology Room
31:$R\leftarrow\emptyset$
32:for $test_{m}$ in $IE$ do
33:     $R_{m}\leftarrow\text{Radiologist}(test_{m},PA)$
34:     $R.\text{append}(R_{m})$
35:end for
36:
37:Back: Consultation Room
38:$FD,TP\leftarrow\emptyset,\emptyset$
39:for $CL_{i}$ in $CT$ do
40:     $FD_{i},TP_{i}\leftarrow CL_{i}(CC,H,V,A,R)$
41:     $FD.\text{append}(FD_{i});\leavevmode\nobreak\ \leavevmode\nobreak\ TP.\text{append}(TP_{i})$
42:end for
43:$FD,TP\leftarrow\text{Aggregate}(FD,CL_{Chair}),\text{Aggregate}(TP,CL_{Chair})$
44:
45:Entry: Treatment Room
46:$\text{Nurse}(PA,FD,TP)$      % Administer treatment
47:
48:Entry: Drug Room
49:Pharmacist($\text{Pharmacist}(PA,FD,TP)$      % Dispense medication
A.3 Evaluation Methods for ClinicalAgent
@@
We continue to use the automatic metrics mentioned in Section 4.2 to evaluate the performance of ClinicalAgent, including DWR, CDR, acceptability, and Avg.
@@
We randomly select one data sample for each disease, totaling 150 cases. We use GPT-4o to score the diagnostic results from seven models mentioned in Section A.3, following the same rubics as human evaluation. The prompt templates used are provided in Table 13 and 14.
A.4 Patient Navigator Settings
Following the experimental design in Section 5.2, we use the Acc@K metric to evaluate navigation accuracy, with K values set at 1, 3, and 5. Additionally, following the description in Section 4.2, we calculate the DIFR-Q and DIFR-N to evaluate instruction following during navigation.
InternLM2, Gemini-Pro, and GPT-4 exhibit outstanding accuracy. Although most models are nearly perfect in the DIFR-Q, they perform poorly in the DIFR-N, especially GPT-4, having a DIFR-N of only 74.03%. Overall, Gemini-Pro and Yi sexhibit remarkable performance. However, Gemini-Pro still has deficiencies in the correctness of department names. We expect the patient navigator to not only accurately guide patients to the appropriate departments but also to minimize errors in department names. Therefore, we choose Yi-Chat as the patient navigator for ClinicalAgent.
@@ 3. Differential Diagnosis:
Prompt for Department Guide (Chinese) {CJK} UTF8gkai 您是一位经验丰富的医院导诊员,根据患者提供的性别、年龄和症状等信息,请从医院科室列表中选择一个患者最需要前往治疗的科室。
请确保您的回答只包含选中的一个科室,不要包含多余的内容。
以下是患者提供的信息:
{Patient Basic Information & Chief Complaint}
以下是医院科室列表:
{Hospital Departments}
您选中的科室:
Prompt for Department Guide (English) You are an experienced hospital guide. Based on the information provided by the patient, including gender, age, and symptoms, please select a department from the list of hospital departments that the patient needs to visit for treatment most.
Please ensure that your response includes only one selected department without any additional content.
The following is the patient’s information:
{Patient Basic Information & Chief Complaint}
The following is the list of hospital departments:
{Hospital Departments}
Your selected department:
Table 10: Prompt templates for department guide.
Prompt for Clinical Diagnosis (Chinese) {CJK} UTF8gkai 您是一位经验丰富的{Department}临床医生,根据给定的病例摘要,请分析并提供一个专业、详细、全面的临床诊断,包含以下6个部分:
1. 初步诊断:分条列出患者可能存在的一种或多种疾病的名称。
2. 诊断依据:分条列出您做出初步诊断的依据。
3. 鉴别诊断:分条列出几种可能导致患者目前症状的疾病,并简要说明为何将它们排除。如果您认为不需要进行鉴别诊断,请直接回答“诊断明确,无需鉴别。”。
4. 主要诊断:一种对患者身体健康危害最大、最需要治疗的疾病的名称。
5. 治疗原则:概述治疗您给出的主要诊断的一般原则。
6. 治疗计划:分条列出初步的治疗计划。
请确保您的回答符合以下格式:
“初步诊断:<您提供的初步诊断>
诊断依据:<您提供的诊断依据>
鉴别诊断:<您提供的鉴别诊断>
主要诊断:<您提供的主要诊断>
治疗原则:<您提供的治疗原则>
治疗计划:<您提供的治疗计划>”
以下是给定的病例摘要:
{Case Summary}
您提供的临床诊断:
Prompt for Clinical Diagnosis (English) You are an experienced {Department} clinician. Based on the given clinical case summary, please analyze and provide a professional, detailed, and comprehensive clinical diagnosis, including the following 6 parts:
1. Preliminary Diagnosis: List the names of one or multiple diseases that the patient might have.
2. Diagnostic Basis: List the basis for your preliminary diagnosis.
3. Differential Diagnosis: List several diseases that could cause the patient’s current symptoms and briefly explain why you exclude them. If you believe differential diagnosis is unnecessary, please directly response “The diagnosis is clear and no differentiation is needed.”
4. Principal Diagnosis: The name of a disease that is most harmful to the patient’s physical health and needs immediate treatment.
5. Therapeutic Principle: Outline the general principles for treating your principal diagnosis.
6. Treatment Plan: List the rough treatment plan.
Please ensure that your response follows this format:
“Preliminary Diagnosis: <Your Preliminary Diagnosis>
Diagnostic Basis: <Your Diagnostic Basis>
Differential Diagnosis: <Your Differential Diagnosis>
Principal Diagnosis: <Your Principal Diagnosis>
Therapeutic Principle: <Your Therapeutic Principle>
Treatment Plan: <Your Treatment Plan>”
The following is the given clinical case summary:
{Case Summary}
Your clinical diagnosis:
Table 11: Prompt templates for clinical diagnosis.
Prompt for Imaging Diagnosis (Chinese) {CJK} UTF8gkai 您是一位经验丰富的放射科医生,根据给定的{Imageological Examination Type}检查报告中的影像所见部分,请分析并分条列出专业的影像诊断。
请确保您的回答高度简洁。
以下是给定的影像所见:
{Findings}
您提供的影像诊断:
Prompt for Imaging Diagnosis (English) You are an experienced radiologist. Based on the findings section of the given {Imageological Examination Type} examination report, please analyze and list professional impression.
Please ensure that your response is highly concise.
The following is the given findings:
{Findings}
Your impression:
Table 12: Prompt templates for imaging diagnosis.
Prompt for GPT-4o Evaluation (Chinese) {CJK} UTF8gkai 您是一位AI评估员,专门评估其他语言模型提供的回答的质量。您的主要目标是根据回答的流畅性、相关性、完整性和医学知识专业性来对回答进行评分。请使用以下评分标准来评估每个指标:
流畅性:
1分:完全破碎且无法阅读的句子片段
2分:大部分破碎,只有少量可读的词
3分:有一定流利度,但词汇有限
4分:在表达复杂主题方面基本上是流畅的
5分:人类水平流利度
相关性:
1分:与问题完全无关
2分:与问题有一定关系,但主要是离题的
3分:相关,但缺乏重点或关键细节
4分:高度相关,解决了大部分问题
5分:直接相关,准确地针对了问题
完整性:
1分:极度不完整
2分:几乎不完整,信息有限
3分:有一定的完整性,有一些信息
4分:大部分信息都已显示
5分:所有信息都已呈现
医学知识专业性:
1分:使用简单明了的语言,没有医学术语
2分:具有一些医学知识,但缺乏深入细节
3分:清晰地传达了一定的复杂医学信息
4分:对医学术语有扎实的认识,但有些细节错误
5分:在所有呈现的医学知识上都是完全正确的
您将会得到以下信息:
- 一个问题
- 这个问题的参考答案
- 一个模型对这个问题给出的回答
[问题]
{Question For Medical Task}
[问题结束]
[参考答案]
{Solution By Human Doctor}
[参考答案结束]
[模型回答]
{Answer By Model}
[模型回答结束]
请确保以JSON格式提供您的的评估结果,并且仅提供JSON,分别对每个提到的指标进行评分,如下例所示:
{“流畅性”: 3, “相关性”: 3, “完整性”: 3, “专业性”: 3}
Table 13: The prompt template for evaluation using GPT-4o.
Prompt for GPT-4o Evaluation (English) You are an AI evaluator specializing in assessing the quality of answers provided by other language models. Your primary goal is to rate the answers based on their fluency, relevance, completeness, proficiency in medicine. Use the following scales to evaluate each criterion:
Fluency:
1: Completely broken and unreadable sentence pieces
2: Mostly broken with few readable tokens
3: Moderately fluent but with limited vocabulary
4: Mostly coherent in expressing complex subjects
5: Human-level fluency
Relevance:
1: Completely unrelated to the question
2: Some relation to the question, but mostly off-topic
3: Relevant, but lacking focus or key details
4: Highly relevant, addressing the main aspects of the question
5: Directly relevant and precisely targeted to the question
Completeness:
1: Extremely incomplete
2: Almost incomplete with limited information
3: Moderate completeness with some information
4: Mostly complete with most of the information displayed
5: Fully complete with all information presented
Proficiency in medicine:
1: Using plain languages with no medical terminology
2: Equipped with some medical knowledge but lacking in-depth details
3: Conveying moderately complex medical information with clarity
4: Showing solid grasp of medical terminology but having some minor mistakes in detail
5: Fully correct in all presented medical knowledge
You will be provided with the following information:
- a question
- the solution to the question
- a model’s answer to the question
[question]
{Question For Medical Task}
[end of question]
[solution]
{Solution By Human Doctor}
[end of solution]
[answer]
{Answer By Model}
[end of answer]
Make sure to provide your evaluation results in JSON format and ONLY the JSON, with separate ratings for each of the mentioned criteria as in the following example:
{“fluency”: 3, “relevance”: 3, “completeness”: 3, “proficiency”: 3}
Table 14: The prompt template for evaluation using GPT-4o. (Cont. 13)

### v2

@@ A.3 ClinicalAgent Algorithm
Departmental Guide: When patient $\mathcal{PA}$ enters the hospital, they submit a simple chief complaint $\mathcal{CC}$ to the patient navigator agent supported by Yi-Chat [37]. The navigator agent, based on the $\mathcal{CC}$ and prior knowledge $\mathcal{PK}$ from Table 2, dynamically arranges the most relevant $\mathcal{K}$ departments $\mathcal{DP}_{\mathit{k}}(k=1,2,\dots,\mathcal{K})$ according to the complexity of the $\mathcal{CC}$, and assigns $\mathcal{N}$ top clinicians $\mathcal{CL}_{\mathit{n}}^{\mathcal{DP}_{\mathit{k}}}(n=1,2,\dots,\mathcal{N})$ from each department $\mathcal{DP}_{\mathit{k}}$ to form the clinician team $\mathcal{CT}$. The $\mathcal{CT}$ prepares for an initial consultation with $\mathcal{PA}$ in the consultation room. The purpose of this stage is to quickly and accurately identify the patient’s initial needs and arrange the appropriate specialized medical team for diagnosis.
[…]
Laboratory Examination: Based on the preliminary consultation advice, the $\mathcal{PA}$ proceeds to the biochemical testing room to undergo a series of necessary tests according to the $\mathcal{LTS}$, such as blood tests and urine analysis. Then, based on the laboratory reports $\mathcal{LR}_{\mathit{i}}$ generated by medical instruments, the biochemist provides the corresponding laboratory diagnostic results $\mathcal{LDR}_{\mathit{i}}$. Finally, the biochemist consolidates all $\mathcal{LDR}_{\mathit{i}}$ to form the final laboratory diagnostic result $\mathcal{LDR}$ and delivers it to the $\mathcal{PA}$.
Imageological Examination: Based on the preliminary consultation advice, the $\mathcal{PA}$ proceeds to the radiology room to undergo a series of necessary examinations according to the imaging test requirements $\mathcal{ITS}$, such as X-rays, CT scans, or MRI. Then, the radiologist provides the corresponding imaging diagnostic results $\mathcal{IDR}_{\mathit{i}}$ based on the textual reports $\mathcal{IR}_{\mathit{i}}$ of the medical images. Finally, the radiologist consolidates all $\mathcal{IDR}_{\mathit{i}}$ to form the final imaging diagnostic result $\mathcal{IDR}$ and delivers it to the $\mathcal{PA}$.
[…]
Medical Treatment: The $\mathcal{PA}$ receives corresponding treatment based on $\mathcal{F}$ and $\mathcal{T}$, including treatments conducted in the treatment room and medication dispensed in the pharmacy. Finally, the $\mathcal{PA}$ leaves the hospital after completing the treatment.
@@
1:Input: Chief Complaint $\mathcal{CC}$, Medical History $\mathcal{MH}$, physical examination $\mathcal{PE}$, Priori Knowledge $\mathcal{PK}$
2:Output: Department of the visit $\mathcal{DP}$, Preliminary Diagnosis $\mathcal{P}$, Diagnostic Basis $\mathcal{B}$, Differential Diagnosis $\mathcal{D}$, Final Diagnosis $\mathcal{F}$, Principle of Treatment $\mathcal{G}$, Treatment Plan $\mathcal{T}$
@@
5:$\mathcal{CT}\leftarrow\emptyset$  % Clinician Team
6:$\mathcal{DP,K,N}$leftarrow Navigator($\mathcal{CC}$)  % $\mathcal{K}$ is the number of arranged departments, $\mathcal{N}$ is the number of clinicians assigned to each department
7:for $k=1$ to $\mathcal{K}$ do
8:    for $n=1$ to $\mathcal{N}$ do
9:         $\mathcal{CL}_{\mathit{n}}^{\mathcal{DP}_{\mathit{k}}}\leftarrow\text{Mapping}(\mathcal{DP}_{\mathit{k}},\mathcal{PK},n)$      % Retrieve and call clinicians
10:         $\mathcal{CT}.\text{append}(\mathcal{CL}_{\mathit{n}}^{\mathcal{DP}_{\mathit{k}}})$
@@
[…]
16:$\mathcal{P},\mathcal{LTS},\mathcal{ITS}\leftarrow\emptyset,\emptyset,\emptyset$  % $\mathcal{P}$ is a set of possible diseases, $\mathcal{LTS}$ is a set of potential laboratory tests that may be required, $\mathcal{ITS}$ is a set of potential imageological tests that may be required
17:for $\mathcal{CL}_{\mathit{i}}\leavevmode\nobreak\ \text{in}\leavevmode\nobreak\ \mathcal{CT}$ do
18:    $\mathcal{P}_{\mathit{i}},\mathcal{LTS}_{\mathit{i}},\mathcal{ITS}_{\mathit{i}}\leftarrow\mathcal{CL}_{\mathit{i}}(\mathcal{CC},\mathcal{MH},\mathcal{PE})$
19:    $\mathcal{P}.\text{append}(\mathcal{P}_{\mathit{i}});\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{LTS}.\text{append}(\mathcal{LTS}_{\mathit{i}});\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{ITS}.\text{append}(\mathcal{ITS}_{\mathit{i}})$
@@
[…]
@@
24:$\mathcal{LDR}\leftarrow\emptyset$  % Laboratory diagnostic results
25:for $test_{j}$ in $\mathcal{LTS}$ do
26:    $\mathcal{LR}_{\mathit{j}}\leftarrow test_{j}(\mathcal{PA})$  % $\mathcal{LR}_{\mathit{j}}$ is a laboratory report generated by medical instruments.
27:    $\mathcal{LDR}_{\mathit{j}}\leftarrow\text{Biochemist}(\mathcal{LR}_{\mathit{j}})$
28:    $\mathcal{LDR}.\text{append}(\mathcal{LDR}_{\mathit{j}})$
29:end for
30:
31:Entry: Radiology Room
32:$\mathcal{IDR}\leftarrow\emptyset$  % Imaging diagnostic results
33:for $test_{m}$ in $\mathcal{ITS}$ do
34:    $\mathcal{IR}_{\mathit{m}}\leftarrow test_{m}(\mathcal{PA})$  % $\mathcal{IR}_{\mathit{m}}$ is a natural language radiological report of medical images.
35:    $\mathcal{IDR}_{\mathit{m}}\leftarrow\text{Radiologist}(\mathcal{IDR}_{\mathit{m}})$
36:    $\mathcal{IDR}.\text{append}(\mathcal{IDR}_{\mathit{m}})$
37:end for
38:
39:Back: Consultation Room
40:$\mathcal{B},\mathcal{D},\mathcal{F},\mathcal{G},\mathcal{T}\leftarrow\emptyset,\emptyset,\emptyset,\emptyset,\emptyset$
41:for $\mathcal{CL}_{\mathit{i}}\leavevmode\nobreak\ \text{in}\leavevmode\nobreak\ \mathcal{CT}$ do
42:    $\mathcal{B}_{\mathit{i}}\leftarrow\mathcal{CL}_{\mathit{i}}(\mathcal{CC},\mathcal{MH},\mathcal{PE},\mathcal{P},\mathcal{LDR},\mathcal{IDR})$
43:    $\mathcal{D}_{\mathit{i}}\leftarrow\mathcal{CL}_{\mathit{i}}(\mathcal{CC},\mathcal{MH},\mathcal{PE},\mathcal{P},\mathcal{LDR},\mathcal{IDR},\mathcal{B}_{\mathit{i}})$
44:    $\mathcal{F}_{\mathit{i}}\leftarrow\mathcal{CL}_{\mathit{i}}(\mathcal{CC},\mathcal{MH},\mathcal{PE},\mathcal{P},\mathcal{LDR},\mathcal{IDR},\mathcal{B}_{\mathit{i}},\mathcal{D}_{\mathit{i}})$
45:    $\mathcal{G}_{\mathit{i}}\leftarrow\mathcal{CL}_{\mathit{i}}(\mathcal{F}_{\mathit{i}})$
46:    $\mathcal{T}_{\mathit{i}}\leftarrow\mathcal{CL}_{\mathit{i}}(\mathcal{CC},\mathcal{MH},\mathcal{PE},\mathcal{P},\mathcal{LDR},\mathcal{IDR},\mathcal{B}_{\mathit{i}},\mathcal{D}_{\mathit{i}},\mathcal{F}_{\mathit{i}},\mathcal{G}_{\mathit{i}})$
47:    .$\mathcal{B}.\text{append}(\mathcal{B}_{\mathit{i}})$; $\mathcal{D}.\text{append}(\mathcal{D}_{\mathit{i}})$; .$\mathcal{F}.\text{append}(\mathcal{F}_{\mathit{i}})$; $\mathcal{G}.\text{append}(\mathcal{G}_{\mathit{i}})$; $\mathcal{T}.\text{append}(\mathcal{T}_{\mathit{i}})$;
48:end for
[…]
51:
52:Entry: Treatment Room
53:$\text{Nurse}(\mathcal{PA},\mathcal{F},\mathcal{T})$      % Administer treatment
54:
55:Entry: Drug Room
56:Pharmacist($\text{Pharmacist}(\mathcal{PA},\mathcal{F},\mathcal{T})$      % Dispense medication
A.4 Evaluation Methods for ClinicalAgent
@@
We continue to use the automatic metrics mentioned in Section 4.2 to evaluate the performance of ClinicalAgent, including $\mathsf{DWR}$, $\mathsf{CDR}$, $\mathsf{Acceptability}$, and $\mathsf{Avg.}$
@@
We randomly select one data sample for each disease, totaling 150 cases. We use GPT-4o to score the diagnostic results from seven models mentioned in Section A.4, following the same rubics as human evaluation.
A.5 Patient Navigator Settings
Following the experimental design in Section 5.2, we use the Acc@K metric to evaluate navigation accuracy, with K values set at 1, 3, and 5. Additionally, following the description in Section 4.2, we calculate the $\mathsf{DIFR\text{-}Q}$ and $\mathsf{DIFR\text{-}N}$ to evaluate instruction following during navigation.
InternLM2, Gemini-Pro, and GPT-4 exhibit outstanding accuracy. Although most models are nearly perfect in the $\mathsf{DIFR\text{-}Q}$, they perform poorly in the $\mathsf{DIFR\text{-}N}$, especially GPT-4, having a $\mathsf{DIFR\text{-}N}$ of only 74.03%. Overall, Gemini-Pro and Yi sexhibit remarkable performance. However, Gemini-Pro still has deficiencies in the correctness of department names. We expect the patient navigator to not only accurately guide patients to the appropriate departments but also to minimize errors in department names. Therefore, we choose Yi-Chat as the patient navigator for ClinicalAgent.